# Differentially Private Matchings

### v2

@@ 6.1 Bounded Arboricity Sparsifiers
Let $S_{G}$ and $S_{G^{\prime}}$ be the sparsified graphs of $G$ and $G^{\prime}$, respectively. Suppose without loss of generality that $G^{\prime}$ contains $E_{\operatorname{extra}}$ additional edges incident to vertex $v$ and $v$ has degree 00 in graph $G$. Then, for each edge $\{v,w\}\in E_{\operatorname{extra}}$, let $e^{w}_{\operatorname{last}}$ be the edge adjacent to $w$ in $G$ whose index in $\pi$ is the last among the edges incident to $w$ in $S_{G}$. If $i_{\pi}(e^{w}_{\operatorname{last}})>i_{\pi}(\{v,w\})$ (where $i_{\pi}(\{v,w\})$ is the index of edge $\{v,w\}$ in $\pi$), then $\{v,w\}$ replaces edge $e^{w}_{\operatorname{last}}$. Since both $G$ and $G^{\prime}$ are simple graphs, at most one edge incident to $w$ gets replaced by an edge in $E_{\operatorname{extra}}$ in $S_{G^{\prime}}$. This set of edge replacements leads to an edge edit distance of $2\Lambda$. ∎
@@
6.2 Proof of Theorem 1.8
6.2.1 Node-DP Maximum Matching with Public Arboricity Bound
In this section, we design a node-DP algorithm to output an implicit matching in the central model using the sparsification techniques derived in Section 6.1 and any edge DP algorithm (e.g. Theorem 1.3) that outputs implicit solutions in the $\varepsilon$-DP setting.
@@ 6.2.1 Node-DP Maximum Matching with Public Arboricity Bound
The algorithm first discards edges according to Theorem 6.2 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 6.1). We can thus run any $(\frac{\varepsilon}{2\Lambda})$-edge DP algorithm to ensure $\varepsilon$-node DP. ∎
6.2.2 Removing the Assumption of Public Bound
@@
In practice, it is not always possible to obtain a good public bound on the arboricity. However, we can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\log(n)/\varepsilon)$ additive error since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a slight blow-up in $b$ and accounting for this blowup leads to a proof of Theorem 1.8.
@@ 7 Improved Node-DP Implicit Bipartite Matchings
In this section, we restate and prove Theorem 1.10. See 1.10
@@
1 Input: Graph $G=(V,E)$, approximation factor $\eta\in(0,1)$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log(n)/(\eta^{4}\varepsilon))$.
@@ 7.1.2 Utility Proof
Combining our two estimates in Eqs. 1 and 2, we have
@@ 7.1.3 Putting it Together
8 Arbitrary Edge-Order Continual Release Implicit Matchings
We now give algorithms for matchings in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. Specifically, we consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: 1) arbitrary-order edge arrival streams and 2) adjacency-list order streams.
In this section, we focus on arbitrary-order edge arrival streams, where edge insertions arrive in an arbitrary order, and show that our edge and node-DP implicit matching algorithms can be implemented in this model. Our results for the second stream type (adjacency-list order streams) is presented in Section 9. To avoid redundancy, we present the edge-DP algorithm (Theorem 1.11), which is essentially a lazy-update version of our static algorithm (Algorithm 2), and then briefly discuss the minor changes needed to obtain a node-DP algorithm (Theorem 1.12). Our results uses a variant of the standard SVT tool that allows us to answer “above” $c$ times, which we recall in Section A.3 for completeness.
See 1.11
Implementing the node-DP arboricity sparsifier of Section 6.2.1 in the continual release model leads to Theorem 1.12, our main continual release node-DP result restated below.
See 1.12
8.1 Proof of Theorem 1.11
We begin with the pseudocode of the $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately in Sections 8.1.1 and 8.1.2, respectively. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.

### v3

@@ 7.1.1 Bounded Arboricity Matching Sparsifier
Let $S_{G}$ and $S_{G^{\prime}}$ be the sparsified graphs of $G$ and $G^{\prime}$, respectively. Suppose without loss of generality that $G^{\prime}$ contains $E_{\operatorname{extra}}$ additional edges incident to vertex $v$ and $v$ has degree 0 in graph $G$. Then, for each edge $\{v,w\}\in E_{\operatorname{extra}}$, let $e^{w}_{\operatorname{last}}$ be the edge adjacent to $w$ in $G$ whose index in $\pi$ is the last among the edges incident to $w$ in $S_{G}$. If $i_{\pi}(e^{w}_{\operatorname{last}})>i_{\pi}(\{v,w\})$ (where $i_{\pi}(\{v,w\})$ is the index of edge $\{v,w\}$ in $\pi$), then $\{v,w\}$ replaces edge $e^{w}_{\operatorname{last}}$. Since both $G$ and $G^{\prime}$ are simple graphs, at most one edge incident to $w$ gets replaced by an edge in $E_{\operatorname{extra}}$ in $S_{G^{\prime}}$. This set of edge replacements leads to an edge edit distance of $2\Lambda$. ∎
@@
7.1.2 Algorithm
In this section, we design a node-DP algorithm to output an implicit matching in the central model assuming a given public bound $\tilde{\alpha}$ on the arboricity $\alpha$ using the sparsification techniques derived in Section 7.1.1 and any edge DP algorithm (e.g. Theorem 2.2) that outputs implicit solutions in the $\varepsilon$-DP setting.
@@ 7.1.2 Algorithm
The algorithm first discards edges according to Theorem 7.3 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 7.2). We can thus run our $(\frac{\varepsilon}{2\Lambda})$-edge DP Algorithm 2 (Theorem 2.2) to ensure $\varepsilon$-node DP. ∎
Public Arboricity Bound.
@@
In practice, it is not always possible to obtain a good public bound on the arboricity. However, we can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\log(n)/\varepsilon)$ additive error since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a slight blow-up in $b$ and accounting for this blowup leads to a proof of Corollary 7.1.
7.2 Node-DP Implicit $b$-Matching (Theorem 2.6)
In this section, we design a node-DP $b$-matching algorithm via arboricity sparsification.
7.2.1 Bounded Arboricity $b$-Matching Sparsifier
$b$-Matching Sparsifier.
Our $b$-Matching sparsifier is identical to our matching sparsifier given in Section 7.1.1, with the exception that the degree threshold $\Lambda$ is adjusted as below.
Theorem 7.5.
$\mu_{b}(H)\leq\mu_{b}(G)\leq(1+\eta)\mu_{b}(H).$
Let $\Lambda_{b}$ be as in Theorem 7.5, $T(v)\subseteq V$ denote some subset of neighbors of $v$ with size at least $\min(\deg_{G}(v),\Lambda_{b})$ for each $v\in V$, and $E_{H}\coloneqq\{uv\in E(G):u\in T(v),v\in T(u)\}$. Then [solomon2021local] showed that $E_{H}$ contains a large $b$-matching for the case of $b=1$. We extend the result to arbitrary $1\leq b\leq n$.
Proof.
We argue by induction on $b$. The base case of $b=1$ is guaranteed by Theorem 7.3. Suppose inductively the statement holds for b1b-1 for some $b\geq 2$ so that our goal is to extend it to the case $b$.
Remark that $M_{1}^{\prime}\cup M_{2}\cup\dots\cup M_{b}$ is a $b$-matching within $G$ by construction. Consider $G_{-1}\coloneqq G-M_{1}^{\prime}$, so that $M_{2}\cup\dots\cup M_{b}$ is a (b1)(b-1)-matching within $G_{-1}$. We claim that $E_{H}\cap E(G_{-1})$ contains a (b1)(b-1)-matching of size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. To prove this claim, we will rely on the induction hypothesis. Define
Then each $T_{-1}(v)$ contains at least $\min(d_{G_{-1}}(v),\Lambda_{b-1})$ neighbors of $v$ in $G_{-1}$. Furthermore, $E_{-1}\subseteq(E_{H}\setminus M_{1}^{\prime})$ by construction. Thus by the induction hypothesis, $E_{-1}$ and hence $H$ contains a $(1+\eta)$-approximate maximum (b1)(b-1)-matching NN of $G_{-1}$.
We have identified disjoint edge subsets $M_{1}^{\prime},N\subseteq E_{H}$ such that $M_{1}^{\prime}$ is a matching in $G$ with size at least $\frac{1}{1+\eta}|M_{1}|$ and $N$ is a (b1)(b-1)-matching in $G-M_{1}^{\prime}$ with size at least $\frac{1}{1+\eta}|M_{2}\cup\dots\cup M_{b}|$. Thus $M_{1}^{\prime}\cup N$ is a $b$-matching in $G$ with size at least
7.2.2 Algorithm
Similar to Section 7, we can implement our implicit $b$-matching algorithm in node-DP setting with the help of the $b$-matching arboricity sparsifier we designed in Section 7.2.1. Then, repeating the steps in Section 7.1 yields a proof of Theorem 2.6.
7.3 Node-DP Implicit Vertex Cover (Theorem 2.7)
In this section, we design a node-DP vertex cover algorithm via arboricity sparsification.
7.3.1 Bounded Arboricity Vertex Cover Sparsifier
Towards proving our node-DP vertex cover algorithm (Theorem 2.7), we describe the arboricity sparsifier for vertex cover.
Let $G_{\leq}^{\Lambda}$ denote the induced subgraph on all vertices with degree less than or equal to some bound $\Lambda$ (which we will set) and $V_{>}^{\Lambda}$ denotes the set of vertices with degree greater than $\Lambda$. [solomon2021local] showed that the union of any $\beta$-approximate vertex cover for $G_{\leq}^{\Lambda}$ and the set of vertices $V_{>}^{\Lambda}$ yields a $(\beta+\eta)$-approximate vertex cover for $G$ for any constant $\eta>0$. We extend the ideas of [solomon2021local] to more general sparsifiers for vertex cover. Our argument holds in greater generality at the cost of an additional additive error of $\operatorname{OPT}(G)$ in the approximation guarantees.
Vertex Cover Sparsifier.
Our vertex cover sparsifier is identical to our matching sparsifier given in Section 7.1.1. Specifically, each vertex keeps the first $\min(\deg_{G}(v),\Lambda)$ incident edges with respect to some fixed ordering $\pi$.
Theorem 7.6.
Let $\eta>0$, $G$ be a graph of arboricity at most $\alpha$, and set
$\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha.$
Let $H\preceq G$ be any subgraph of $G$ such that $G_{\leq}^{\Lambda}\preceq H\preceq G$. Here $\preceq$ denotes the subgraph relation. If $C$ is a $(\beta,\zeta)$-approximate vertex cover of $H$, then $C\cup V_{>}^{\Lambda}$ is a cover of $G$ with cardinality at most
$(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$
Lemma 7.7 (Lemma 3.8 in [solomon2021local]).
Let $\eta>0$ and $G$ be a graph of arboricity at most $\alpha$ and set $\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha$. If $C$ is any vertex cover for $G$, and $U_{>}:=V_{>}\setminus C$, then $|U_{>}|\leq\eta|C|$.
Remark 7.8.
Lemma 3.8 in [solomon2021local] holds as long as any induced subgraph of $G$ has average degree at most $2\alpha$, which is certainly the case for graphs of arboricity at most $\alpha$. Moreover, [solomon2021local] stated their result for
$V_{\geq}=\{v\in V:\deg(v)\geq\Lambda\},\qquad U_{\geq}=V_{\geq}\setminus C.$
But since $U_{>}\subseteq U_{\geq}$, the result clearly still holds. In order to remain consistent with our matching sparsifier, we state their result with $V_{>}$ and $U_{>}$.
Proof of Theorem 7.6.
Any edge in $H$ is covered by $C$, which is a $(\beta,\zeta)$-approximate vertex cover of $H$, and any edge that is not in $H$ is covered by $V_{>}^{\Lambda}$. Thus $C\cup V_{>}^{\Lambda}$ is indeed a vertex cover. By Lemma 7.7, $|V_{>}^{\Lambda}\setminus C^{*}|\leq\eta|C^{*}|$ where $C^{*}$ is a minimum vertex cover of $G$. Then,
$|C\cup V_{>}^{\Lambda}|$ $\leq|C|+|V_{>}^{\Lambda}|$ (1)
$\leq\beta\operatorname{OPT}(H)+\zeta+|V_{>}^{\Lambda}\cap C^{*}|+|V_{>}^{\Lambda}\setminus C^{*}|$ (2)
$\leq\beta\operatorname{OPT}(G)+\zeta+\operatorname{OPT}(G)+\eta\operatorname{OPT}(G)$ (3)
$=(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$ (4)
Eq. 3 follows by Lemma 7.7 and since $|V_{>}^{\Lambda}\cap C^{*}|\leq\operatorname{OPT}(G)$ due to $C^{*}$ being an optimal cover for $G$. ∎
Let $H$ denote our vertex cover sparsifier. Remark that we are guaranteed to have $G_{\leq}^{\Lambda}\subseteq H$ since if an edge $\{u,w\}$ has $\deg(u)\leq\Lambda$ and $\deg(w)\leq\Lambda$, then the edge is guaranteed to be in the first $\Lambda$ edges of the ordering $\pi$ for both of its endpoints.
7.3.2 Algorithm
We now present our node-DP vertex cover algorithm that releases an implicit vertex cover (Definition 7.11) using the sparsification techniques derived in Section 7.3.1. As a subroutine, we use the static edge DP implicit vertex cover algorithm (Theorem 7.10) of [GuptaLMRT10]. Similar to our node-DP matching algorithm, we first describe an algorithm assuming a public bound $\tilde{\alpha}$ on the arboricity $\alpha$.
Theorem 7.9.
Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit vertex cover (Definition 7.11). Moreover, if $\tilde{\alpha}\geq\alpha$, the implicit solution is a $O(\nicefrac{{\tilde{\alpha}}}{{\varepsilon}})$-approximation with probability $0.99$.
Before proving Theorem 7.9, we first define the precise implicit solution released by our algorithm.
Value vs Solution.
The node-DP algorithm to estimate the value of a maximum matching immediately yields a node-DP algorithm for estimating the size of a minimum vertex cover. In fact, it suffices to compute the size of a greedy maximal matching, which has a coupled global sensitivity of 1 with respect to node neighboring graphs. This is then a $2$-approximation for the size of a minimum vertex cover. However, it may be desirable to output actual solutions to the vertex cover problem.
[GuptaLMRT10] demonstrated how to output an implicit solution through an ordering of the vertices. Specifically, the theorem below returns an implicit solution where every edge is considered an entity and can determine, using the released implicit solution, which of its endpoints covers it.
Theorem 7.10 (Theorem 5.1, Theorem 5.2 in [GuptaLMRT10]).
Fix $\beta\in(0,1)$. There is a polynomial time $\varepsilon$-edge DP algorithm EdgePrivateVC that outputs an implicit vertex cover with expected cardinality at most
$\left(2+\frac{16}{\varepsilon}\right)\operatorname{OPT}.$
By Markov’s inequality, running EdgePrivateVC guarantees a $O(\nicefrac{{1}}{{\varepsilon}})$-approximate solution with probability $0.999$.
Implicit Vertex Cover.
The implicit vertex cover from [GuptaLMRT10] is an ordering of the vertices where each edge is covered by the earlier vertex in the ordering. Such an implicit ordering allows for each edge to determine which of its endpoints covers it given the public ordering.
In our setting, if we run EdgePrivateVC on a sparsified graph $H$, the ordering we output is with respect to $H$ and some edges that were deleted from $G$ to obtain $H$ may not be covered by the intended solution. Thus we also output the threshold we input into the sparsifier so that an edge is first covered by a high-degree endpoint if it has one, and otherwise by the earlier vertex in the ordering.
Definition 7.11 (Implicit Vertex Cover).
An implicit solution is a pair
$\left(\pi,\Lambda\right)\in S_{n}\times\mathbb{Z}_{\geq 0}$
where $\pi$ is an ordering of the vertices and $\Lambda$ is the threshold used to perform the sparsification. Every edge is first covered by an endpoint with degree strictly greater than $\Lambda$ if it exists, and otherwise by the endpoint which appears earlier in $\pi$.
Remark 7.12 (Ordering-Only Implicit Vertex Cover).
We assume each edge is an entity that knows its endpoints as well as their (private) degrees. Ideally, we would utilize the exact same solution concept as [GuptaLMRT10], i.e., output an ordering only, which does not require each edge to have access to the degrees of its endpoints for decoding. However, we prove that any $(\varepsilon,\delta)$-node DP algorithm which outputs a [GuptaLMRT10] implicit solution necessarily incurs large $\Omega(n)$-error (Section 7.5).
With the precise notion of an implicit vertex cover in hand, we are now ready to prove Theorem 7.9.
Proof Sketch.
[…]
Public Arboricity Bound.
We can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\nicefrac{{1}}{{\varepsilon}})$ additive error with probability $0.999$ since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a constant factor in the approximation ratio and leads to a proof of Theorem 2.7.
7.4 Lower Bound for Releasing Sparsifiers
See 2.5
Proof.
Fix $V$ with $|V|=n$, and let $r\in V$. For any $S\subseteq V\setminus\{r\}$, let $G_{S}$ be the graph whose edge set consists of an edge from $r$ to all nodes in $S$ (i.e., a star from $r$ to $S$ and all other nodes having degree 0). Note that all such graphs have arboricity $1$. A similar symmetry argument to the other lower bounds implies that without loss of generality, we may assume that the probability that $\mathcal{A}$ includes some edge $\{r,v\}$ is the same for all $v\in S$ (we will denote this probability by $p_{S}$).
Fix some node $v\in S\setminus\{r\}$, and consider what happens when we run $\mathcal{A}$ on the graph $G_{\{v\}}$. Since $\mathbb{E}[\nu(E^{\prime}\cap E)]\geq\frac{\nu(G_{\{v\}})}{1+\eta}$ by assumption, we know that $p_{\{v\}}=\Pr[\{r,v\}\in E^{\prime}]\geq\frac{1}{1+\eta}$.
7.5 Lower Bound for Ordering-Only Node-DP Implicit Vertex Cover
Let $G=(V,E)$ be a graph. As mentioned in Section 7.3, the implicit vertex cover used by [GuptaLMRT10] is an ordering. More formally, given an ordering $\pi$ of $V$, we let $G_{\pi}$ be the directed graph where for any edge $\{u,v\}\in E$ we direct it $(v,u)$ if $\pi(u)<\pi(v)$ and otherwise we direct it $(u,v)$. In other words, we direct it into the endpoint that is earlier in the ordering. Given $\pi$, let $S(\pi,G)=\{v\in V:\exists(u,v)\in E(G_{\pi})\}$. Then clearly $S(\pi,G)$ is a vertex cover of $G$, which we think of as the vertex cover defined implicitly by $\pi$.
Theorem 7.13.
Let $\mathcal{A}$ be an algorithm which satisfies $(\epsilon,\delta)$-node DP and outputs an ordering $\pi$ of $V$. Then the expected approximation ratio of $\mathcal{A}$ is at least $\frac{n}{4}e^{-\epsilon}\left(1-2\delta\right)$.
Proof.
Let $V$ be a set of $n$ nodes. For each $v\in V$, let $G_{v}$ be the star graph with vertex set $V$ with center node $v$. Let $G_{\emptyset}$ denote the empty graph with vertex set $V$. Note that $G_{\emptyset}$ and $G_{v}$ are neighboring graphs in the node DP setting for all $v\in V$.
Consider the distribution over orderings generated by running $\mathcal{A}$ on $G_{\emptyset}$. For every $v\in V$, let $p_{v}=\Pr_{\pi\sim\mathcal{A}(G_{\emptyset})}[\pi(v)\geq n/2]$ denote the probability that $v$ is in the second half of the nodes in the ordering. Then clearly $\sum_{v\in V}p_{v}=n/2$, since the expected number of nodes in the second half is exactly $n/2$ and is also (by linearity of expectations) equal to $\sum_{v\in V}p_{v}$. Thus there exists some node $v\in V$ such that $p_{v}\geq 1/2$.
Now consider $G_{v}$. Let $p^{\prime}_{v}=\Pr_{\pi\sim\mathcal{A}(G_{v})}[\pi(v)\leq n/2]$ be the same probability but when $\mathcal{A}$ is run on $G_{v}$. Then by node differential privacy, we get that $p_{v}\leq e^{\epsilon}p^{\prime}_{v}+\delta$, and hence $p^{\prime}_{v}\geq e^{-\epsilon}\left(p_{v}-\delta\right)\geq e^{-\epsilon}\left(\frac{1}{2}-\delta\right)$. Note that if $v$ is in the second half of $\pi$, then $|S(\pi,G_{v})|\geq n/2$, since every node in the first half of $\pi$ would be in $S(\pi,G_{v})$. Thus we have that
$\mathbb{E}_{\pi\sim\mathcal{A}(G_{v})}[|S(\pi,G_{v})|]$ $\geq\frac{n}{2}p^{\prime}_{v}\geq\frac{n}{2}e^{-\epsilon}\left(\frac{1}{2}-\delta\right)=\frac{n}{4}e^{-\epsilon}\left(1-2\delta\right).$
On the other hand, it is obvious that the optimal vertex cover in $G_{v}$ has size $1$ (it is just $\{v\}$). This implies the theorem. ∎
@@ 8 Improved Node-DP Implicit Bipartite Matchings
In this section, we restate and prove Theorem 2.8. See 2.8
8.1 Detailed Algorithm Description
@@
1Input: Graph $G=(V,E)$, approximation factor $\eta\in(0,1)$, privacy parameter $\varepsilon>0$, matching parameter $b=\Omega(\log(n)/(\eta^{4}\varepsilon))$.
@@ 8.2.2 Utility Proof
Combining our two estimates in Eqs. 5 and 6, we have
@@ 8.2.3 Putting it Together
9 Continual Release of Implicit Matchings
We now give algorithms for matchings in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. Specifically, we consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: 1) arbitrary-order edge arrival streams under both edge and node-DP and 2) adjacency-list order streams.
Edge-DP for Arbitrary Edge Order Streams.
We first focus on arbitrary-order edge arrival streams, where edge insertions arrive in an arbitrary order, and show that our edge and node-DP implicit matching algorithms can be implemented in this model. To avoid redundancy, we present the edge-DP algorithm (Theorem 2.9), which is essentially a lazy-update version of our static algorithm (Algorithm 2), and then briefly discuss the minor changes needed to obtain a node-DP algorithm (Theorem 2.10).
See 2.9
As a special case, we recover the following result for matchings.
Corollary 9.1.
For $\varrho\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with probability $1-\frac{1}{n^{c}}$, each $b$-matching is a $\left(2+\varrho,O\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximation with respect to the maximum 1-matching size.
Similar to Theorem 5.2, we can remove the need for a bicriteria approximation when the degree bound $b$ is slightly higher.
Theorem 9.2.
For $\eta\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching is a $\left(2+\eta,\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $b$-matching.
Node-DP for Arbitrary Edge Order Streams.
Implementing the node-DP arboricity sparsifier of Section 7.1 in the continual release model leads to Theorem 2.10, our main continual release node-DP result restated below.
See 2.10 Similar to the edge-DP case, the case of $b^{\prime}=1$ leads to a special case with approximation guarantees with respect to the maximum 1-matching size.
Corollary 9.3.
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and $b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}\right)$. There is an $\varepsilon$-node DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching has the size of a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $1$-matching.
Adjacency-List Order Streams.
In the adjacency-list order insertion streams, nodes arrive in an arbitrary order, and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. Remark that in the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream. See 2.11
Organization.
In Section 9.1, we present our $\varepsilon$-edge DP arbitrary edge-order continual release algorithm as well as the corollaries in Section 9.1.3. Then, we achieve stronger $\varepsilon$-node DP guarantees in Section 9.2. Finally, we present the details of our adjacency-list continual release algorithm in Section 9.3.
9.1 Edge-DP Arbitrary Edge-Order Continual Release Implicit $b$-Matchings
We begin with the pseudocode of the $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately in Sections 9.1.1 and 9.1.2, respectively. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.
Our results use a variant of the standard SVT tool that allows us to answer “above” $c$ times, which we recall in Section A.3 for completeness.